OpenStudy (anonymous):

If the bases are 1/2 and 1/4, and the height is 1, what is the area? 1/8 sq. units 3/8 sq. units 3/4 sq. units

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

Its a trapezoid

ganeshie8 (ganeshie8):

use the area formula : http://0.tqn.com/d/math/1/S/4/G/trapezoidr.gif

OpenStudy (anonymous):

I got the equation. A=1/2 (1) (1/2+1/4)

ganeshie8 (ganeshie8):

a = 1/2 b = 1/4 h = 1 plug in

ganeshie8 (ganeshie8):

and you're right !

OpenStudy (anonymous):

So I would get0.375

OpenStudy (anonymous):

Ty for help and medal

ganeshie8 (ganeshie8):

leave it in fractions

OpenStudy (anonymous):

3/8

ganeshie8 (ganeshie8):

1/2 + 1/4 what do get when u add a `quarter` to a `half` ?

ganeshie8 (ganeshie8):

yep ! it is 3/8

OpenStudy (anonymous):

Alright thanks so much

ganeshie8 (ganeshie8):

yw!